Ndou Drilling company has total cost function is TC=25Q2+30Q-55, and their marginal cost function is MC= 50Q+30, with demand of P=150-15Q, and whose marginal revenue function MR= 150-30Q, where Q is output and P is price.

Assume that the firm maximizes profit but cannot practice price discrimination



How much does the firm charge?
1
Expert's answer
2021-07-08T14:32:24-0400

"MC= 50Q+30"

"MR= 150-30Q"

Then, set MR = MC to solve for Q:

"50Q+30= 150-30Q"

"50Q+30Q= 150-30"

"80Q= 120"

"Q=1.5"

When Q = 1.5, then "P=150-15Q"

"P=150-15*1.5"

P= $"127.50"